8- You are planning to save for retirement over the next 15 years. To do this, you will invest $1,100 a month in a stock account and $500 a month in a bond account. The return on the stock account is expected to be 7 percent, and the bond account will pay 4 percent. When you retire, you will combine your money into an account with a 5 percent return. How much can you withdraw each month during retirement assuming a 20-year withdrawal period?
Answer #1

step 1, find the future value of your deposits

Find fv of stocks,  pmt=-1100, i=7/12, n=15*12=180, and compute fv=348658.53 as below

FV = PV[(1+r)^n - 1] / r

= 1100 [(1+0.005833333)^180 - 1 ]/ 0.005833333

= 348658.5264

Find fv of bond,  pmt=-500, i=4/12, n=15*12=180, and compute fv=123045.24

FV = PV[(1+r)^n - 1] / r

= 500 [(1+0.003333333)^180 - 1 ]/ 0.003333333

= 123045.2441

This gives a total of 471703.77.

Use this fv as the new pv in 15 years and solve for payment.  pv=471703.77, n=20*12=240, i=5/12, fv=0, and compute the pmt = 3113.04

PV = P [(1 - (1+r)^-n] / r

471703.77 = P [(1 - (1+0.004167)^ -240] / 0.004167

471703.77 = P * 151.5253

P = 471703.77 / 151.5253

P = 3113.036 i.e. 3113.04